OpenStudy (anonymous):

already in factored form. How would you solve (4x+1)(x-2) = 0?

Parth (parthkohli):

If the product of two numbers is zero, then one of them must be zero.

Parth (parthkohli):

So 4x + 1 or x-2=0

OpenStudy (potatoes.ramu):

x=-1/4 or x=2...

OpenStudy (cggurumanjunath):

4x+1=0 x-2=0

OpenStudy (anonymous):

\[x=\frac{ -1 }{ 4}\] or \[x=2\]